What are the magnitude Δυ and angle θ of the change in velocity if the initial velocity is 50.0 m/s south and the final veloci is 25.0 m/s west? Let the angle θ be measured relative to the east direction (e.g., θ 90 means Δυ points north, θ-180. means Δυ points west, etc.).
Answer #1

Considering cartesian coordinates; i.e, east, west, north south are +x, -x, +y, -y axes respectively

vi = -50 j ; vf = -25 i;

change in velocity, dv = vf - vi = -25i + 50j

its magnitude, = sqrt(252 + 502) = 55.90 m/s

angle, measured clockwise from east direction,= tan-1(50/25) = 63.434o

as measured anticlockwise

\theta = 360 - 63.434 = 296.566o